OpenStudy (anonymous):

I really need help How would I go about solving: (x^4/3 * x^7/3)/(x^2/3), (^7sqrt(x))^21, x^1/3 * x^2/3 * x^4/3 * ^3sqrt(x)

OpenStudy (anonymous):

Take a picture of it and attach it.

OpenStudy (anonymous):

\[\frac{ x^{\frac{ 4 }{ 3 }} * x^{\frac{ 7 }{ 3 }} }{x ^{\frac{ 2 }{ 3 }} }\]

OpenStudy (anonymous):

Thats the first one

OpenStudy (anonymous):

I got something that will help

OpenStudy (anonymous):

Ok

OpenStudy (anonymous):

I tagged you in something

OpenStudy (anonymous):

So the product would be x^28/9 right?